OpenStudy (anonymous):

Identify the curve by finding the Cartesian Equation: r = 4sec(theta) so the formulas that can be used here are: x = rcos(theta) y = rsin(theta) x^2 + y^2 = r^2 tan(theta) = y/x I need to get it to be in the form x^2 + y^2 = r^2 I did: r = 4sec(theta) r = 4(r/x) x = 4 but now... I'm stuck because no matter what I do with that x = 4, I'm don't end up with x^2 + y^2 = (number) format. Any help would be appreciated. Ty.

myininaya (myininaya):

\[r=4 \frac{r}{x} =>rx=4r => x=4\] That is right.

myininaya (myininaya):

If we had \[r=4\cos(\theta)\] \[r=4 \frac{x}{r} => r^2=4x => x^2+y^2=4x \]

myininaya (myininaya):

or even \[r=4\sin(\theta) \] \[r=4 \cdot \frac{y}{r} => r^2=4y => x^2+y^2=4y \]

OpenStudy (anonymous):

not sure if I am correct. \[r ^{2}=16\sec ^{2}\theta =16\left( 1+\tan ^{2}\theta \right)=16\left( 1+\frac{ y ^{2} }{x ^{2} } \right)\] \[x ^{2}+y ^{2}=16\left( \frac{ x ^{2}+y ^{2} }{x ^{2} } \right)\] solve it.

OpenStudy (anonymous):

But once I get x=4, how do i convert that into an partesian equation?

myininaya (myininaya):

x=4 is just a vertical line going through x=4 on the x-axis.